# Global Dynamics of McKean-Vlasov SDEs via Stochastic Order

### v1

@@ 2 Assumptions and Main Results
(i) The continuous function $V^{\prime}:\mathbb{R}\to\mathbb{R}$ is one-sided Lipschitz, i.e., there exists $L>0$ such that, for all $x,y\in\mathbb{R}$,
$-(x-y)(V^{\prime}(x)-V^{\prime}(y))\leq L\left\lvert x-y\right\rvert^{2}.$
@@
$-(x-y)(V^{\prime}(x)-V^{\prime}(y))\leq-\alpha\left\lvert x-y\right\rvert^{2+\delta}+\beta.$
@@
The stochastic order “$\leq_{\mathrm{st}}$” on $\mathcal{P}_{2}(\mathbb{R})$ is a partial order, defined by
$\mu\leq_{\mathrm{st}}\nu\text{ if and only if }\int_{\mathbb{R}}f\mathop{}\!\mathrm{d}\mu\leq\int_{\mathbb{R}}f\mathop{}\!\mathrm{d}\nu\text{ for all bounded increasing functions $f\colon\mathbb{R}\to\mathbb{R}$}.$
Definition 2.2.
Definition 2.3.
  1. (i)
    $-2xV^{\prime}(x+a)-2x(W^{\prime}*\mu)(x)+\sigma^{2}(x+a)\leq-g(x^{2},\|\mu\|_{2}^{2});$
  2. (ii)
    the number $r>0$ and the function $g$ satisfy
@@
    $\mathcal{W}_{2}(\nu_{t},\nu)\to 0\text{ as $t\to-\infty$ and }\mathcal{W}_{2}(\nu_{t},\tilde{\nu})\to 0\text{ as $t\to\infty$}.$
@@
Now we state our main theorems. Theorem 2.1, 2.3 and Corollary 2.2, 2.4 are abstract results for the equation (1.1), while Theorem 2.5-2.9 deal with concrete equations. In the sequel, we denote $P_{t}^{*}:\mathcal{P}_{2}(\mathbb{R})\to\mathcal{P}_{2}(\mathbb{R})$ by the semigroup generated by the equation (1.1), see also (4.5).
@@
Suppose Assumption 1 holds, and suppose the zero-crossing number $Z(V^{\prime})$ of $V^{\prime}:\mathbb{R}\to\mathbb{R}$ is $(2n-1)$. If the equation (1.1) is locally dissipative at $a_{1}<a_{2}<\cdots<a_{n}$ with configurations $\{(r_{k},g_{k})\}_{k=1}^{n}$, and if
$r_{k}^{2}+r_{k+1}^{2}\leq\frac{1}{2}(a_{k+1}-a_{k})^{2}\ \text{ for all $k=1,2,\dots,n-1$},$ (2.1)
@@
    (Total Orderedness) there are exactly $(2n-1)$ invariant measures $\{\nu_{k}\}_{k=1}^{2n-1}$ with $\nu_{1}<_{\mathrm{st}}\nu_{2}<_{\mathrm{st}}\cdots<_{\mathrm{st}}\nu_{2n-1}$;
@@
(ii). By definition, a connecting orbit $\{\nu_{t}\}_{t\in\mathbb{R}}$ from $\nu$ to $\tilde{\nu}$ as a bounded set itself cannot be attracted by its forward limit $\tilde{\nu}$, i.e., $\mathcal{W}_{2}(P_{t}^{*}\{\nu_{s}:s\in\mathbb{R}\},\tilde{\nu})\nrightarrow 0$, although $\mathcal{W}_{2}(P_{t}^{*}\nu_{s},\tilde{\nu})\to 0$ as $t\to\infty$ for any $s\in\mathbb{R}$. So, the set of invariant measures cannot attract all bounded subsets of $\mathcal{P}_{2}(\mathbb{R})$. In other word, the order interval $[\underline{\nu},\overline{\nu}]$ in Theorem 2.1 (ii) cannot be replaced by the set of invariant measures.
(iii). By the alternating arrangement of the invariant measures in Theorem 2.3, we have more information on the basins of attraction (see Theorem 4.2).
  • if $\mu<_{\mathrm{st}}\nu^{1,\downarrow}_{t_{0}}$ for some $t_{0}\in\mathbb{R}$, then $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{1})\to 0$ as $t\to\infty$;
  • if $\nu^{k-1,\uparrow}_{t_{0}}<_{\mathrm{st}}\mu<_{\mathrm{st}}\nu^{k,\downarrow}_{t_{0}}$ for some $t_{0}\in\mathbb{R}$, $k=2,3,\dots,n-1$, then $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{2k-1})\to 0$ as $t\to\infty$;
  • if $\mu>_{\mathrm{st}}\nu^{n-1,\uparrow}_{t_{0}}$ for some $t_{0}\in\mathbb{R}$, then $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{2n-1})\to 0$ as $t\to\infty$.
@@
(i). For each $k=1,2,\dots,2n-2$, although the set $\{\mu_{m}\}_{m\in(m_{k},m_{k+1})}$ is a totally ordered arc with endpoints $\nu_{k}$, $\nu_{k+1}$, it is not a connecting orbit between invariant measures. One can check it from the Fokker-Planck equation.
(ii). Corollary 2.4 also provides more information on basins of attraction. For each $k=1,2,\dots,n$, if $\mu_{m}\leq_{\mathrm{st}}\mu\leq_{\mathrm{st}}\mu_{\tilde{m}}$ for some $m_{2k-2}<m<\tilde{m}<m_{2k}$, then by the order-preservation of $P^{*}$ and the squeezing principle, we have $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{2k-1})\to 0$ as $t\to\infty$.
Besides abstract results, we have the following applications.
@@
$\theta\geq\frac{27(9+\sqrt{17})}{128},\quad 0<\underline{\sigma}^{2}<\sigma^{2}(x)<\frac{51\sqrt{17}-107}{256}\ \text{ for all $x\in\mathbb{R}$},$ (2.5)
and let $r=\frac{9-\sqrt{17}}{8}$, then the following statements are true.
@@
Theorem 2.6 (Double-well with flat bottoms, Figure 1).
Consider the following one-dimensional McKean-Vlasov SDE,
where
$\theta\geq\frac{27(9+\sqrt{17})}{128},\quad 0<\underline{\sigma}^{2}<\sigma^{2}(x)<\frac{51\sqrt{17}-171}{256}\ \text{ for all $x\in\mathbb{R}$},$ (2.7)
Theorem 2.7 (Asymmetric double-well, Figure 1).
Consider the following one-dimensional McKean-Vlasov SDE,
$\mathop{}\!\mathrm{d}X_{t}=-\left[X_{t}(X_{t}-1)(X_{t}+2)+\theta\left(X_{t}-\mathbf{E}[X_{t}]\right)\right]\mathop{}\!\mathrm{d}t+\sigma(X_{t})\mathop{}\!\mathrm{d}B_{t}.$ (2.8)
$\theta\geq\frac{4(3+\sqrt{3})}{3},\quad 0<\underline{\sigma}^{2}<\sigma^{2}(x)<\frac{3(39+55\sqrt{33})}{256}\ \text{ for all $x\in\mathbb{R}$},$
and let $r_{-2}=\frac{(\sqrt{33}-15)^{2}}{64}$, $r_{1}=\frac{6-3\sqrt{3}}{2}$, then the following statements are true.
  1. (i)
    There are exactly three invariant measures, $\nu_{-2},\nu_{0},\nu_{1}\in\mathcal{P}_{\infty}(\mathbb{R})$ with $\nu_{-2}<_{\mathrm{st}}\nu_{0}<_{\mathrm{st}}\nu_{1}$;
  2. (ii)
    For all bounded sets $B\subset\mathcal{P}_{2}(\mathbb{R})$, $\mathcal{W}_{2}(P_{t}^{*}B,[\nu_{-2},\nu_{1}])\to 0$ as $t\to\infty$;
  3. (iii)
    There are a decreasing connecting orbit from $\nu_{0}$ to $\nu_{-2}$ and an increasing connecting orbit from $\nu_{0}$ to $\nu_{1}$;
  4. (iv)
    There hold $\nu_{-2}\in B(\delta_{-2},r_{-2})$, $\nu_{1}\in B(\delta_{1},r_{1})$, and
    $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{-2})\to 0\ \text{ as $t\to\infty$ for all $\mu\in\bigcup_{a\leq 0}B(\delta_{-2+a},r_{-2})$},$
    $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{1})\to 0\ \text{ as $t\to\infty$ for all $\mu\in\bigcup_{a\geq 0}B(\delta_{1+a},r_{1})$}.$
Figure 2: Phase diagram for multi-well landscapes
@@
$\theta\geq 8\sqrt{5+\sqrt{13}},\quad 0<\underline{\sigma}^{2}<\sigma^{2}(x)<\frac{4(13\sqrt{13}-35)}{27}\ \text{ for all $x\in\mathbb{R}$},$
and let $r=\frac{5-\sqrt{13}}{3}$, then the following statements are true.
@@
Theorem 2.9 (Double-well with perturbation).
@@
$\mathop{}\!\mathrm{d}X_{t}=-\left[X_{t}(X_{t}-1)(X_{t}+1)+f(X_{t})+\theta\left(X_{t}-\mathbf{E}[X_{t}]\right)\right]\mathop{}\!\mathrm{d}t+\sigma(X_{t})\mathop{}\!\mathrm{d}B_{t}.$ (2.10)
If the parameter $\theta$ and the Lipschitz functions $f\colon\mathbb{R}\to\mathbb{R}$, $\sigma\colon\mathbb{R}\to\mathbb{R}$ satisfy
$\theta\geq 0,\quad\left\lvert f(x)\right\rvert\leq L,\ \ 0<\underline{\sigma}^{2}<\sigma^{2}(x)<\overline{\sigma}^{2}\ \text{ for all $x\in\mathbb{R}$},$
then the exactly same results as in Theorem 2.1 are valid.
@@ 3.1 Preliminaries of stochastic order
Let $\mathcal{P}(\mathbb{R})$ be the set of probability measures on $\mathbb{R}$, and let $\mathcal{P}_{p}(\mathbb{R})$ be the set of those having finite $p$-th moments,
@@
$\mathcal{P}_{p}(\mathbb{R}):=\left\{\mu\in\mathcal{P}(\mathbb{R}):\left\lVert\mu\right\rVert_{p}=\left(\int_{\mathbb{R}}\left\lvert x\right\rvert^{p}\mathop{}\!\mathrm{d}\mu(x)\right)^{1/p}<\infty\right\}.$
@@
For $p\in[1,\infty)$, equip $\mathcal{P}_{p}(\mathbb{R})$ with the $p$-Wasserstein metric defined as below,
$\mathcal{W}_{p}(\mu,\nu):=\inf\left\{\left(\mathbf{E}[\left\lvert X-Y\right\rvert^{p}]\right)^{1/p}:\mathcal{L}(X)=\mu,\,\mathcal{L}(Y)=\nu\right\},$
@@
where $\mathcal{L}(\cdot)$ represents the law of a random variable. Then $(\mathcal{P}_{p}(\mathbb{R}),\mathcal{W}_{p})$ is a complete metric space (see Villani [68, Theorem 6.18]).
Next, we review the so-called “stochastic order”, which is also known as stochastic domination. By Definition 2.1, for $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R})$,
$\mu\leq_{\mathrm{st}}\nu\text{ if and only if }\int_{\mathbb{R}}f\mathop{}\!\mathrm{d}\mu\leq\int_{\mathbb{R}}f\mathop{}\!\mathrm{d}\nu\text{ for all bounded increasing functions $f\colon\mathbb{R}\to\mathbb{R}$}.$
@@
Then “$\leq_{\mathrm{st}}$” is a partial order on $\mathcal{P}_{2}(\mathbb{R})$. Indeed, the reflexivity and transitivity are obvious, and the antisymmetry can be obtained by taking $f=\mathbf{1}_{(x,\infty)}$ for any $x\in\mathbb{R}$. Moreover, as shown in [48, Lemma 3.2], the stochastic order is a closed partial order relation in $\mathcal{P}_{2}(\mathbb{R})$, namely, the subset $\{(\mu,\nu):\mu\leq_{\mathrm{st}}\nu\}$ is closed in the product space $\mathcal{P}_{2}(\mathbb{R})\times\mathcal{P}_{2}(\mathbb{R})$. That is to say, $\mu\leq_{\mathrm{st}}\nu$ whenever $\mathcal{W}_{2}(\mu_{n},\mu)\to 0$ and $\mathcal{W}_{2}(\nu_{n},\nu)\to 0$ as $t\to\infty$ and $\mu_{n}\leq_{\mathrm{st}}\nu_{n}$ for all $n$.
@@
If $\mu\leq_{\mathrm{st}}\nu$ and $\mu\neq\nu$, we write $\mu<_{\mathrm{st}}\nu$. The reversed signs $\geq_{\mathrm{st}}$, $>_{\mathrm{st}}$ are used in the usual way. For $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R})$ with $\mu\leq_{\mathrm{st}}\nu$, the order interval is defined by $[\mu,\nu]=\{\rho\in\mathcal{P}_{2}(\mathbb{R}):\mu\leq_{\mathrm{st}}\rho\leq_{\mathrm{st}}\nu\}$, $(\mu,\nu]=\{\rho\in\mathcal{P}_{2}(\mathbb{R}):\mu<_{\mathrm{st}}\rho\leq_{\mathrm{st}}\nu\}$ and similarly we can define $[\mu,\nu)$ and $(\mu,\nu)$. Measures $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R})$ are said to be order related, if $\mu\leq_{\mathrm{st}}\nu$ or $\nu\leq_{\mathrm{st}}\mu$ holds. Otherwise, they are order unrelated. Given two subsets $A$ and $B$ of $\mathcal{P}_{2}(\mathbb{R})$, we write $A\leq_{\mathrm{st}}B$ ($A<_{\mathrm{st}}B$) when $\mu\leq_{\mathrm{st}}\nu$ ($\mu<_{\mathrm{st}}\nu$) holds for each choice of $\mu\in A$ and $\nu\in B$. Sets $A,B\subset\mathcal{P}_{2}(\mathbb{R})$ are said to be order related, if $A\leq_{\mathrm{st}}B$ or $B\leq_{\mathrm{st}}A$ holds. A totally ordered set (which is also referred as chain) $A$ in $\mathcal{P}_{2}(\mathbb{R})$ means that, any two measures $\mu$ and $\nu$ in $A$ are order related.
@@
A measure $\mu\in\mathcal{P}_{2}(\mathbb{R})$ is an upper bound (lower bound) of a set $A\subset\mathcal{P}_{2}(\mathbb{R})$ if $\mu\geq_{\mathrm{st}}A$ ($\mu\leq_{\mathrm{st}}A$). If the set $A$ has an upper (lower) bound, it is said to be bounded from above (below). A set $A$ which is bounded from both above and below is said to be order bounded. An upper bound $\mu_{0}$ of $A$ is said to be the supremum of $A$, denoted by $\mu_{0}=\sup_{\leq_{\mathrm{st}}}A$, if any other upper bound $\mu$ satisfies $\mu\geq_{\mathrm{st}}\mu_{0}$. Similarly, a lower bound $\mu_{0}$ of $A$ is said to be the infimum of $A$, denoted by $\mu_{0}=\inf_{\leq_{\mathrm{st}}}A$, if any other lower bound $\mu$ satisfies $\mu\leq_{\mathrm{st}}\mu_{0}$. By the antisymmetry of the stochastic order, if the supremum (infimum) exists, it must be unique.
@@
There are some complementary material of the stochastic order that we will use later. Firstly, a useful equivalent characterization of the stochastic order is Strassen’s theorem (see e.g., Lindvall [47, equation (3)]), which says $\mu\leq_{\mathrm{st}}\nu$ if and only if there exist two random variables $X,Y$ such that
@@
Secondly, according to Hiai-Lawson-Lim [40, Proposition 3.11], test functions $f$ in Definition 2.1 can be restricted as bounded continuous increasing. If $\mu$, $\nu$ both have finite first moments, it is equivalent for test functions ranging over all $1$-Lipschitz increasing functions (see Fritz-Perrone [36, Theorem 4.2.1]), where a $1$-Lipschitz function $f$ means $\left\lvert f(x)-f(y)\right\rvert\leq\left\lvert x-y\right\rvert$ for all $x,y\in\mathbb{R}$. We may confine test functions in such different sets whenever convenient. Additionally, an important observation regarding the definition of $\mu\leq_{\mathrm{st}}\nu$, is that when we take the test function as $f=-\mathbf{1}_{(-\infty,x]}$ or $f=\mathbf{1}_{(x,\infty)}$, we have $\mu\big{(}(-\infty,x]\big{)}\geq\nu\big{(}(-\infty,x]\big{)}$ and $\mu\big{(}(x,\infty)\big{)}\leq\nu\big{(}(x,\infty)\big{)}$, for any $x\in\mathbb{R}$.

### v2

@@ 1.1 Assumptions and Main Results
(i) The function $V\in C^{2}(\mathbb{R};\mathbb{R})$ such that the continuous function $V^{\prime}:\mathbb{R}\to\mathbb{R}$ is one-sided Lipschitz, i.e., there exists $L>0$ such that, for all $x,y\in\mathbb{R}$,
$(x-y)(V^{\prime}(x)-V^{\prime}(y))\leq L\left\lvert x-y\right\rvert^{2}.$
@@
$(x-y)(V^{\prime}(x)-V^{\prime}(y))\leq-\alpha\left\lvert x-y\right\rvert^{2+\delta}+\beta.$
@@
The stochastic order $\leq_{\mathrm{st}}$” on $\mathcal{P}_{2}(\mathbb{R})$ is a partial order, defined by
[…]
@@
[…]
@@
Now we state our main theorems. Theorem 1.2, 1.4 and Corollary 1.3, 1.5 are abstract results for the equation (1.1), while Theorem 1.6-1.8 deal with concrete equations. In the sequel, we denote by $P_{t}^{*}:\mathcal{P}_{2}(\mathbb{R})\to\mathcal{P}_{2}(\mathbb{R})$ the semigroup associated to the equation (1.1):
$P_{t}^{*}\mu:=\mathcal{L}(X_{t}^{\xi})\ \text{ with }\ \mathcal{L}(\xi)=\mu,$ (1.4)
where $X_{t}^{\xi}$ is the unique solution of (1.1) with initial data $\xi$ at time 0.
@@
Definition 1.3.
Definition 1.4.
  1. (i)
    $2xV^{\prime}(x+a)-2x(W^{\prime}*\mu)(x)+\sigma^{2}(x+a)\leq-g(x^{2},\|\mu\|_{2}^{2});$
  2. (ii)
    the number $r>0$ and the function $g$ satisfy
Remark 1.2.
It is straightforward to verify that Definition 1.4 provides an equivalent formulation of [36, Definition 1.2] in the one-dimensional setting. The concept of local dissipation was introduced by Zhang [57], and was later applied by Feng-Qu-Zhao [18] to the time-inhomogeneous McKean-Vlasov equation and by Bao-Wang [4] to the McKean–Vlasov SDEs with jumps.
@@
Suppose Assumption 1 holds, and suppose the zero-crossing number $Z(V^{\prime})$ of $V^{\prime}:\mathbb{R}\to\mathbb{R}$ is (2n1)(2n-1). If the equation (1.1) is locally dissipative at $a_{1}<a_{2}<\cdots<a_{n}$ with configurations $\{(r_{k},g_{k})\}_{k=1}^{n}$, and if
$r_{k}+r_{k+1}\leq a_{k+1}-a_{k},\ \text{ for all $k=1,2,\dots,n-1$},$ (1.5)
@@
    (Total Orderedness) there are exactly (2n1)(2n-1) invariant measures $\{\nu_{k}\}_{k=1}^{2n-1}$ with $\nu_{1}<_{\mathrm{st}}\nu_{2}<_{\mathrm{st}}\cdots<_{\mathrm{st}}\nu_{2n-1}$;
@@
(ii). By definition, a connecting orbit $\{\nu_{t}\}_{t\in\mathbb{R}}$ from $\nu$ to $\tilde{\nu}$ as a bounded set itself cannot be attracted by its forward limit $\tilde{\nu}$, i.e., $\mathcal{W}_{2}(P_{t}^{*}\{\nu_{s}:s\in\mathbb{R}\},\tilde{\nu})\nrightarrow 0$, although $\mathcal{W}_{2}(P_{t}^{*}\nu_{s},\tilde{\nu})\to 0$ as $t\to\infty$ for any $s\in\mathbb{R}$. So, the set of invariant measures cannot attract all bounded subsets of $\mathcal{P}_{2}(\mathbb{R})$. In other word, the order interval $\underline{\nu},\overline{\nu}]$ in Theorem 1.2 (ii) cannot be replaced by the set of invariant measures.
We next present a corollary providing the order position of invariant measures and also explicit expressions for certain densities in the basins of attraction. Consider a family of probability measures $\{\mu_{m}\}_{m\in\mathbb{R}}$ on $\mathbb{R}$, whose densities are given by
@@
For each $k=1,2,\dots,2n-2$, although the set $\{\mu_{m}\}_{m\in(m_{k},m_{k+1})}$ is a totally ordered arc with endpoints $\nu_{k}$, $\nu_{k+1}$, it is not a connecting orbit between invariant measures. One can check it from the Fokker-Planck equation.
As an application of our abstract theorems, we present below the classical double-well and multi-well examples. In contrast to Theorem 1.1, the confining potential $V$ is given explicitly in these cases, which allows us to determine explicit parameter ranges and to provide explicit descriptions of attracting sets. On the other hand, Proposition 4.3 provides a detailed analysis of the number of invariant measures under multiplicative noise, thereby verifying the a priori assumption in the main result of [36]. Based on this analysis, the subsequent examples not only address basins of attraction with respect to both the partial order and the topology, but also strengthen the results on connecting orbits for the corresponding examples considered in [36]. In particular, even under multiplicative noise, the classical double-well equation (1.9) admits connecting orbits; for the multi-well equation (1.11) we identify exactly five invariant measures together with a precise heteroclinic structure connecting them via monotone connecting orbits.
@@
$\theta\geq\frac{27(9+\sqrt{17})}{128},\quad 0<\inf_{x\in\mathbb{R}}\sigma^{2}(x)\leq\sup_{x\in\mathbb{R}}\sigma^{2}(x)<\frac{51\sqrt{17}-107}{256},$ (1.10)
and let $r=\frac{9-\sqrt{17}}{8}$, then
@@
$\theta\geq 8\sqrt{5+\sqrt{13}},\quad 0<\inf_{x\in\mathbb{R}}\sigma^{2}(x)\leq\sup_{x\in\mathbb{R}}\sigma^{2}(x)<\frac{4(13\sqrt{13}-35)}{27},$
and let $r=\frac{\sqrt{15-3\sqrt{13}}}{3}$, then
@@
Figure 2: Phase diagram for multi-well landscapes
Owing to our result on the number of invariant measures (Proposition 4.3), our framework is also capable of handling the case where $V^{\prime}$ vanishes on an interval.
Theorem 1.8 (Double-well with flat bottoms, Figure 1).
@@
where
$\theta\geq\frac{27(9+\sqrt{17})}{128},\quad 0<\inf_{x\in\mathbb{R}}\sigma^{2}(x)\leq\sup_{x\in\mathbb{R}}\sigma^{2}(x)<\frac{51\sqrt{17}-171}{256},$ (1.13)
@@ 2.1 Preliminaries of stochastic order
Let $\mathcal{P}(\mathbb{R}^{d})$ be the set of probability measures on $\mathbb{R}^{d}$, and let $\mathcal{P}_{p}(\mathbb{R}^{d})$ be the set of those having finite $p$-th moments,
@@
$\mathcal{P}_{p}(\mathbb{R}^{d}):=\left\{\mu\in\mathcal{P}(\mathbb{R}^{d}):\left\lVert\mu\right\rVert_{p}=\left(\int_{\mathbb{R}^{d}}\left\lvert x\right\rvert^{p}\mathop{}\!\mathrm{d}\mu(x)\right)^{1/p}<\infty\right\}.$
@@
For $p\in[1,\infty)$, equip $\mathcal{P}_{p}(\mathbb{R}^{d})$ with the $p$-Wasserstein metric defined as below,
$\mathcal{W}_{p}(\mu,\nu):=\inf\Big\{\big(\mathbf{E}[\left\lvert X-Y\right\rvert^{p}]\big)^{1/p}:\mathcal{L}(X)=\mu,\,\mathcal{L}(Y)=\nu\Big\},$
@@
where $\mathcal{L}(\cdot)$ represents the law of a random variable. Then $(\mathcal{P}_{p}(\mathbb{R}^{d}),\mathcal{W}_{p})$ is a complete metric space (see Villani [54, Theorem 6.18]).
Next, we review the so-called “stochastic order”, which is also known as stochastic domination; we refer to, e.g., Kamae-Krengel-O’Brien [32], Lindvall [34, Chapter IV] for a detailed discussion. Give $\mathbb{R}^{d}$ an order relation by coordinate-wise comparison, i.e., $x\leq y$ if and only if $x_{i}\leq y_{i}$ for all $i=1,2,\dots,d$. For $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, the stochastic order is defined by the following,
[…]
@@
where “increasing” means $f(x)\leq f(y)$ whenever $x\leq y$. Then $\leq_{\mathrm{st}}$ is a partial order on $\mathcal{P}_{2}(\mathbb{R}^{d})$. Indeed, the reflexivity and transitivity are obvious, and the antisymmetry can be obtained by taking $f=\mathbf{1}_{(x,\infty)^{d}}$ for any $x\in\mathbb{R}^{d}$. Moreover, as shown in [36, Lemma 2.1], the stochastic order is a closed partial order relation in $\mathcal{P}_{2}(\mathbb{R}^{d})$, namely, the subset $\{(\mu,\nu):\mu\leq_{\mathrm{st}}\nu\}$ is closed in the product space $\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathcal{P}_{2}(\mathbb{R}^{d})$. That is to say, $\mu\leq_{\mathrm{st}}\nu$ whenever $\mathcal{W}_{2}(\mu_{n},\mu)\to 0$ and $\mathcal{W}_{2}(\nu_{n},\nu)\to 0$ as $t\to\infty$ and $\mu_{n}\leq_{\mathrm{st}}\nu_{n}$ for all $n$.
@@
If $\mu\leq_{\mathrm{st}}\nu$ and $\mu\neq\nu$, we write $\mu<_{\mathrm{st}}\nu$. The reversed signs $\geq_{\mathrm{st}}$, $_{\mathrm{st}}$ are used in the usual way. For $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ with $\mu\leq_{\mathrm{st}}\nu$, the order interval is defined by $\mu,\nu]=\{\rho\in\mathcal{P}_{2}(\mathbb{R}^{d}):\mu\leq_{\mathrm{st}}\rho\leq_{\mathrm{st}}\nu\}$, $(\mu,\nu]=\{\rho\in\mathcal{P}_{2}(\mathbb{R}^{d}):\mu<_{\mathrm{st}}\rho\leq_{\mathrm{st}}\nu\}$ and similarly we can define $\mu,\nu)$ and $(\mu,\nu)$. Measures $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ are said to be order related, if $\mu\leq_{\mathrm{st}}\nu$ or $\nu\leq_{\mathrm{st}}\mu$ holds. Otherwise, they are order unrelated. Given two subsets $A$ and $B$ of $\mathcal{P}_{2}(\mathbb{R}^{d})$, we write $A\leq_{\mathrm{st}}B$ ($A<_{\mathrm{st}}B$) when $\mu\leq_{\mathrm{st}}\nu$ ($\mu<_{\mathrm{st}}\nu$) holds for each choice of $\mu\in A$ and $\nu\in B$. Sets $A,B\subset\mathcal{P}_{2}(\mathbb{R}^{d})$ are said to be order related, if $A\leq_{\mathrm{st}}B$ or $B\leq_{\mathrm{st}}A$ holds. A totally ordered set (which is also referred as chain) $A$ in $\mathcal{P}_{2}(\mathbb{R}^{d})$ means that, any two measures $\mu$ and $\nu$ in $A$ are order related.
@@
A measure $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ is an upper bound (lower bound) of a set $A\subset\mathcal{P}_{2}(\mathbb{R}^{d})$ if $\mu\geq_{\mathrm{st}}A$ ($\mu\leq_{\mathrm{st}}A$). If the set $A$ has an upper (lower) bound, it is said to be bounded from above (below). A set $A$ which is bounded from both above and below is said to be order bounded. An upper bound $\mu_{0}$ of $A$ is said to be the supremum of $A$, denoted by $\mu_{0}=\sup_{\leq_{\mathrm{st}}}A$, if any other upper bound $\mu$ satisfies $\mu\geq_{\mathrm{st}}\mu_{0}$. Similarly, a lower bound $\mu_{0}$ of $A$ is said to be the infimum of $A$, denoted by $\mu_{0}=\inf_{\leq_{\mathrm{st}}}A$, if any other lower bound $\mu$ satisfies $\mu\leq_{\mathrm{st}}\mu_{0}$. By the antisymmetry of the stochastic order, if the supremum (infimum) exists, it must be unique.
@@
There are some complementary material of the stochastic order that we will use later. Firstly, a useful equivalent characterization of the stochastic order is Strassen’s theorem (see e.g., Lindvall [34, Theorem IV.2.4] or [35, equation (3)]), which says $\mu\leq_{\mathrm{st}}\nu$ if and only if there exist two random variables $X,Y$ such that
@@
Secondly, according to Hiai-Lawson-Lim [25, Proposition 3.11], test functions $f$ in Definition 1.1 can be restricted as bounded continuous increasing. If $\mu$, $\nu$ both have finite first moments, it is equivalent for test functions ranging over all $1$-Lipschitz increasing functions (see Fritz-Perrone [19, Theorem 4.2.1]), where a $1$-Lipschitz function $f$ means $\left\lvert f(x)-f(y)\right\rvert\leq\left\lvert x-y\right\rvert$ for all $x,y\in\mathbb{R}^{d}$. We may confine test functions in such different sets whenever convenient. Additionally, an important observation regarding the definition of $\mu\leq_{\mathrm{st}}\nu$, is that when we take the test function as $f=-\mathbf{1}_{(-\infty,x]^{d}}$ or $f=\mathbf{1}_{(x,\infty)^{d}}$, we have $\mu\big((-\infty,x]^{d}\big)\geq\nu\big((-\infty,x]^{d}\big)$ and $\mu\big((x,\infty)^{d}\big)\leq\nu\big((x,\infty)^{d}\big)$, for any $x\in\mathbb{R}^{d}$.